OpenStudy (anonymous):

horizontal translation 6 units right f(x)=x

OpenStudy (mathmale):

Mind writing that as the original problem? I'd bet the original problem read like this: "Write the function that results when the graph of f(x)=x is translated horiz. 6 units to the right." You may not think so, but being accurate like that may actually help you to do better in solving problems like this one on your own. But anyway. If, for example, g(x) = x^2 - 1, and we translate the graph of g(x) 3 units to the right, the resulting new function will be: h(x) = (x-3)^2 - 1. Please use this example to attempt rewriting your f(x) in modified form to reflect that translation to the right by 6 units.

OpenStudy (anonymous):

This what my book tells me: "Let g(x) be the indicated transformation(s) of f(x)=x. Write the rule for g(x)." Then it gives me the problem. " horizontal translation 6 units right."

OpenStudy (mathmale):

Thanks. More words in the book's version of the problem, but the problem is much clearer because of those extra words. Hope my example (above) helps; mind trying it? Then we can get back to your original question.

OpenStudy (anonymous):

So it would be g(x)= (x-6)?

OpenStudy (mathmale):

Yes. perfect. Magnifique. Excellent.